Status Report

NASA ISS On-Orbit Status 4 July 2009

By SpaceRef Editor
July 4, 2009
Filed under , , ,
NASA ISS On-Orbit Status 4 July 2009

All ISS systems continue to function nominally, except those noted previously or below. U.S. Independence Day.
Saturday – off-duty day for CDR Gennady Padalka (Russia), FE-1 Michael Barratt (USA), FE-2 Koichi Wakata (Japan), FE-3 Roman Romanenko (Russia), FE-4 Robert Thirsk (Canada), FE-5 Frank DeWinne (Belgium).

The ISS crew is back on nominal wake/sleep cycle (2:00am – 5:30pm EDT).

Upon wakeup, FE-1 Barratt, FE-4 Thirsk & FE-5 DeWinne continued their current session of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), logging data from their Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. [To monitor the crewmember’s sleep/wake patterns and light exposure, the crewmembers wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

The six-member crew performed the regular weekly three-hour task of thorough station cleaning, including COL (Columbus Orbital Laboratory) and Kibo. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, Padalka & Romanenko conducted regular maintenance inspection & cleaning on fan screens in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9), plus dust filter replacement in the FGB.

Later, FE-3 Romanenko performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

At ~9:10am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

After temporarily deactivating the ALTEA (Anomalous Long Term Effects on Astronauts) DOSI equipment and swinging it out of the way, FE-4 Thirsk worked on the OGS (Oxygen Generator System) in the Lab, first clearing access to the OGS rack by removing equipment from its right side, then configuring the OGS equipment for activation by the ground, including connecting the H2 sensor’s oxygen QDs (Quick Disconnects). [Afterwards, the equipment was replaced at its stowage location and ALTEA reactivated.]

FE-5 DeWinne retrieved and stowed the four passive FMK (Formaldehyde Monitoring Kit) sampling assemblies deployed by him on 7/2 in the Lab (at P3, below CEVIS) and SM (at the most forward handrail, on panel 307), to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis on the ground. [Two monitors each are usually attached side by side, preferably in an orientation with their faces perpendicular to the direction of air flow.]

FE-1 Barratt performed the periodic status check on the MERLIN (Microgravity Experiment Research Locker Incubator) Galley fridge, replacing the desiccants and looking for any internal condensation moisture. [MERLIN is used for cold storage of crew food and drink.]

Frank completed the periodic recharge of five power tool (Makita) batteries, to be used for ICV (Integrated Cardiovascular) and NIFM (Non-Intrusive Flow Meter) operations.

FE-2 Wakata performed the regular bi-monthly rebooting of the OCA Router and File Server SSC (Station Support Computer) laptops.

Working off the Russian discretionary “time permitting” task list, CDR Padalka was to conduct another run for Russia’s Environmental Safety Agency (EKON), making observations and taking KPT-3 aerial photography of environmental conditions on earth using the Nikon D2X with the SIGMA 300-800mm telephoto lens.

The crew completed their regular daily 2.5-hr. physical workout program on the CEVIS cycle ergometer (FE-1), TVIS treadmill with vibration isolation (CDR, FE-3, FE-4, FE-5), RED resistive exercise device (CDR, FE-1, FE-2, FE-4, FE-5) and VELO cycle ergometer with bungee cord load trainer (FE-2, FE-3). [The interim RED is being used in lieu of the ARED (Advanced RED) until the latter has had its damaged VIS (Vibration Isolation System) dashpot replaced and can be put back in service.]

Later, Mike Barratt transferred the exercise data file to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

Weekly Science Update (Expedition Twenty — Week 4)

3-D SPACE: Third session successfully completed on 6/30. “Thanks Mike!”

AgCam (Agricultural Camera): “Mike: Thank you for your work with troubleshooting AgCam earlier in the week. The PD team is analyzing results of the diagnostics test and the files. New troubleshooting steps will come up as soon as they are ready.”

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): Microscope target test successfully completed on 6/19.

Biological Rhythms (JAXA): No report.

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: “Koichi, Bob, thanks for completing your pill ingestion. Your next session is scheduled for 7/6.”

CARD (Long Term Microgravity: A Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): CB periodic maintenance was completed. Ground has confirmed that it is fully functional.

CBEF (JAXA Cell Biology Experiment Facility): No report.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Bob, your next ICV with CCISS session is scheduled for 7/6. You will be performing the CCISS Baro Study right after the setup, so the Holter will not be near the end of the recording."

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): See MDCA/Flex.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Completed.

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the PCDF (Protein Crystallization Diagnostic Facility) experiment. EDR is providing power / data and temperature control (via cooling loop) to PCDF.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): PCBA S/W Upgrade was successfully completed on 7/2, in preparation for SOLO execution.

EPO (Educational Payload Operations, NASA): “Thank you for completing the EPO-SLAMMD-Demo. Great job! Your ingenuity is greatly appreciated. EPO is happy with the outcome.”

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): Ongoing.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): Complete.

EPO Moon Score (JAXA): Planned.

EPO Try Zero-G (JAXA): “No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): No report.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): EuTEF platform and its instruments have been running nominally during this reporting period. On-going science acquisition for DOSTEL, EXPOSE, MEDET. For FIPEX, a new EOP script will be started on 7/2 and will continue running till 7/6. The DEBIE-2 instrument continues to regularly generate empty science packets, and as a work-around the instrument script includes a daily power cycle command, which is skipped from ground if the science packets remain nominal. After IMMT approval on 6/25, PLEGPAY Experiment 2 could be re-activated and was performed on 6/29. Data are under assessment. Langmuir Probe measurements performed on 7/1 and 7/2.

FACET (JAXA): No report.

FSL (Fluid Science Laboratory): MMA data downlink performed on 7/1. FSL Optical Target #4 check-out was completed by 7/1.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): “Complete downloading the data of Koichi’s Holter experiment which was performed on GMT 141-142 (Kibo Week 7).”

HQPC (JAXA): To be launched by 34P.


ICV (Integrated Cardiovascular): “Bob and Frank: Thanks for another successful resting echo scan. We had a tight window for comm, but you got all the views we needed and the ground team was very pleased with the quality! We are looking forward to next week’s Ambulatory monitoring session. This go around will be a little different from the first: you will NOT erase the Cardiopres memory (the FD14 data will be transferred again along with the new data) and there will be a few extra button presses during the set-up of the Cardiopres to aid in the troubleshooting efforts from the FD14 session. Thanks again for all the crew notes from the first monitoring session as they really helped the ground team track your progress!”

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): “PD team is evaluating the pictures you took during the vial shake and bubble check activity. Thanks for your work.KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators.”

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): “Marangoni Experiment unique equipment set up for 2nd series of Marangoni Experiment (MEIS2) was completed. Very good job. Thank you, Bob, for your work. During this task, Bob reported that the missing LED cap was not found. We will continue the task because the cap itself does not affect the function. Experiment team will continue to monitor when they start the experiment.

MAXI (JAXA): Ongoing.

MDCA/Flex: No report.

Micro-G Clay (JAXA EPO): Complete.

MMA (JAXA/Microgravity Measurement Apparatus): MMA data downlink performed on 7/1. FSL Optical Target #4 check-out was completed by 7/1.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): No report.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): The current cycle is extended to allow the observed crystals to develop further. After another PCDF reboot on 6/27, a new script was started with reactor temperature decreased to +19.5 degC. This increased the growth as expected and did not lead to additional nucleation. On 7/01, it was decided to extend the same run further and to decrease the temperature by another 0.5 degC.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Mike, Bob & Frank, you are currently completing another week of sleep logging this week. We downlinked the data last week and it all looks complete. Thanks for your participation. The next download/initialization session is scheduled for 7/10 for the FE-1, FE-2, FE-4, and FE-5 Actiwatches. The Actiwatches will stop taking data on 7/13.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): The Sun visibility window (#17) ended on 6/26. SOLAR software upgrade was performed on 7/1. A zero procedure was tested successfully to confirm the upgrade. Further confirmation of the S/W upgrade is expected at the start of the next Sun visibility window (expected to start around 7/20). SOLSPEC Calibration performed on 7/1.

SOLO (Sodium Loading in Microgravity): PCBA software upgrade was successfully completed on 7/2, in preparation for SOLO execution.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.

SPICE (Smoke Point In Co-flow Experiment): No report.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): Through 6/30, the ground has received a total of 16,404 of ISS CEO imagery for review and cataloguing. [“We are pleased to confirm your acquisition of views of the following targets: Vredefort Impact Crater, South Africa – excellent context views, as we requested; Maracaibo, Venezuela – excellent context and tight views with little cloudiness – target requirements complete; and Aswan Airport – target acquired – soft focus – will request again. Your skilful, low-light view of Big Thompson Mesa, Utah will be published on NASA/GSFC’s Earth Observatory website this weekend. It beautifully illustrates the geologic structure and sedimentary rock layers of a portion of the Water Pocket Fold in the Colorado Plateau where more than 100 million years of sediment deposition lies exposed in a rugged and remote desert setting. Nice shot!”]

No CEO (Crew Earth Observation) photo targets uplinked for today due to the short workday.

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

Summary of Increment 19 & 20 Research Investigations:

Summary of Increment 19 & 20 Research Investigations:

1. Human Research & Countermeasures

. 3D-Space
. Biorhythms
. Braslet (SDTO)
. Bisphosphonates
. DOSIS (alternate)
. Integrated Cardiovascular
. Integrated Immune
. Neurospat
. Nutrition/Repository
. Portable PFS C/O
. Spinal Elongation
. Thermolab/EKE (Contingency)
. Visual Performance
. VO2max (Contingency)

2. Biological Science

. CERIS (alternate)
. CSI-03 (alternate)
. Microbe-1
. RAD Silk
. Space Seed

3. Physical Science

. BCAT 4
. BCAT 5
. CIR/MDCA/FLEX (alternate)
. FIR/LMM/CVB (alternate)
. FOAM Stability (alternate)
. G1
. Ice Crystal
. InSpace-2 (alternate)
. Marangoni Surface
. PCDF (alternate)

4. Technology Development

. Altea-DOSI
. Area Dosimeter
. Avatar
. CGBA (alternate)
. ENOSE (alternate)
. RaDI-N (alternate)
. SAMS (alternate)
. SpaceDRUMS (alternate)
. VCAM (alternate)

5. Education & Earth Observation

. AgCam
. EarthKAM
. JAXA Commercial 2
. JAXA Commercial 3
. JAXA Commercial 4
. JAXA EPO Forest
. Photo-Moon
. Space Clothes
. Space Poem
. Tomatosphere
. TRY Zero-G

Significant Events Ahead (all dates Eastern Time, some changes possible!):
07/11/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD; (7:39am EDT)
07/12/09 — Progress 33P Re-rendezvous attempt (based on solar constraints)
07/13/09 — STS-127/Endeavour/2J/A docking (if launched nominally 7/11)
07/24/09 — Progress 34P launch
07/25/09 — STS-127/Endeavour/2J/A undocking
07/27/09 — STS-127/Endeavour/2J/A landing (KSC, ~12:16pm EDT)
07/27/09 — Progress 34P docking (if STS-127 departs nominally; can slip to 7/29)
07/31/09 — PMA-3 relocation
08/18/09 — STS-128/Discovery/17A – MPLM (P), LMC (~4:25am EDT)
09/08/09 — H-IIB (JAXA HTV-1) launch
09/16/09 — H-IIB (JAXA HTV-1) berth
09/29/09 — Progress 34P undock
09/30/09 — Soyuz TMA-16/20S launch
10/02/09 — Soyuz TMA-16/20S docking (SM aft, until MRM-2 w/new port)
10/08/09 — H-IIB (JAXA HTV-1) unberth
10/11/09 — Soyuz TMA-14/18S undock
10/15/09 — Progress 35P launch
11/10/09 — 5R/MRM-2 (Russian Mini Research Module 2) on Soyuz-U
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/07/09 — Soyuz TMA-17/21S launch
12/26/09 — Progress 36P launch
02/03/10 — Progress 37P launch
02/04/10 — STS-130/Endeavour/20A – Node-3 + Cupola
03/18/10 — STS-131/Discovery/19A – MPLM(P), LMC
04/02/10 — Soyuz TMA-18/22S launch
04/27/10 — Progress 38P launch
05/14/10 — STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1
05/29/10 — Soyuz TMA-19/23S launch
06/25/10 — Progress 39P launch
07/29/10 — STS-133/Endeavour/ULF5 – ELC4, MPLM
08/11/10 — Progress 40P launch
09/16/10 — STS-134/Discovery/ULF6 – ELC3, AMS
09/29/10 — Soyuz TMA-20/24S launch
10/19/10 — Progress 41P launch
11/??/10 — ATV2 – Ariane 5 (ESA)
12/??/11 — 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton

Note: The daily ISS On-Orbit Status reports can also be found at

SpaceRef staff editor.